OpenStudy (naga447):

Generate the first 5 terms of this sequence: Please help!! t-t f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (naga447):

2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

if n =3, then f(n) is f(3) right? n-1 =? and f(n-1) = f(?) n-2 =? and f(n-2) =f(?)

OpenStudy (naga447):

n - 1 = 2? so f(n-1) = f(2)?

OpenStudy (naga447):

I don't understand what this means could you elaborate?

OpenStudy (naga447):

@OOOPS

OpenStudy (anonymous):

answer each of them n =3, n-1=??

OpenStudy (naga447):

if n = 3 then it would be 2 but what do I need to do after this to get the other unknown terms

OpenStudy (naga447):

f(3) = f(3-1) + f(3-2)

OpenStudy (anonymous):

yup, 3-1=2, right? and 3-2 =1, right? so it becomes f(3) = f(2) +f(1) , you have f(1) , f(2) already, just plug in

OpenStudy (anonymous):

got what I meant??

OpenStudy (anonymous):

f(3-1) =f(2) right??

OpenStudy (naga447):

yeah

OpenStudy (anonymous):

so, f(3)=??

OpenStudy (naga447):

f(3) = f(2+1)?

OpenStudy (anonymous):

and given f(2) = 3 and f(1)=2 so, just plug the values in \[f(3) = f(3-1)+f(3-2)\\f(3) = f~~(2) ~~~~+~f(1)\\f(3) = ~~~~~3~~~~~~+~~~2 ~~=~~5\] Got what I meant??

OpenStudy (naga447):

Yeah I think I understand f(3) = 5

OpenStudy (anonymous):